\begin{thebibliography}{4}%文字数が最大の文献番号を書く

\bibitem{AraNori}
D. Arapura and M. Nori.
\textit{Solvable fundamental groups of algebraic varieties and Kähler manifolds.}
Compositio Mathematica \textbf{116} (1999), no.~2, 173--188.

\bibitem{RodCam}
R. Aguilar and F. Campana.
\textit{The nilpotent quotients of quasi-projective manifolds with proper quasi-Albanese map.}
\texttt{arXiv:2301.11232v2} (2023).

\bibitem{Cam}
F. Campana.
\textit{Remarques sur les groupes de Kähler nilpotents.}
Annales Scientifiques de l’École Normale Supérieure (Série 4) \textbf{28} (1995), no.~3, 307--316.

\bibitem{Cam2}
F. Campana.
\textit{Orbifolds, special varieties and classification theory.}
Annales de l’Institut Fourier (Grenoble) \textbf{54} (2004), no.~3, 499--630.

\bibitem{CadDengYam}
B. Cadorel, Y. Deng, and K. Yamanoi.
\textit{Hyperbolicity and fundamental groups of complex quasi-projective varieties.}
\texttt{arXiv:2212.12225} (2022).

\bibitem{Del}
P. Deligne.
\textit{Théorie de Hodge. II.}
Publications Mathématiques de l’Institut des Hautes Études Scientifiques \textbf{40} (1971), 5--57.

\bibitem{Malcev}
A. I. Malcev.
\textit{On a class of homogeneous spaces.}
Izvestiya Akademii Nauk SSSR. Seriya Matematicheskaya \textbf{13} (1949), 9--32.

\bibitem{Morgan}
J. W. Morgan.
\textit{The algebraic topology of smooth algebraic varieties.}
Publications Mathématiques de l’Institut des Hautes Études Scientifiques \textbf{48} (1978), 137--204.

\bibitem{Nil}
M. Goze and Y. Khakimdjanov.
\textit{Nilpotent Lie Algebras.}
Mathematics and its Applications, vol.~361,
Kluwer Academic Publishers, Dordrecht, (1996).

\bibitem{PeterSteen}
C. A. M. Peters and J. H. M. Steenbrink.
\textit{Mixed Hodge Structures.}
Ergebnisse der Mathematik und ihrer Grenzgebiete. 3. Folge, vol.~52,
Springer-Verlag, Berlin–Heidelberg, (2008).

\bibitem{Pierre}
P. Pierre.
\textit{Lectures on K\" ahler Groups.}
Princeton Mathematical Series (2025).

\bibitem{Raghu}
M. S. Raghunathan.
\textit{Discrete Subgroups of Lie Groups.}
Ergebnisse der Mathematik und ihrer Grenzgebiete, Band~68,
Springer-Verlag, Berlin–Heidelberg, (1972).

\bibitem{Rogov}
V. Rogov.
\textit{O-minimal geometry of higher Albanese manifolds.}
\texttt{arXiv:2505.07632} (2025).

\bibitem{grading}
T. Shimoji.
\textit{Gradings on nilpotent Lie algebras associated with the nilpotent fundamental groups of smooth complex algebraic varieties.}
\texttt{arXiv:2504.08571} (2025).

\bibitem{JohVen}
A. J. Sommese and A. Van de Ven.
\textit{Homotopy groups of pullbacks of varieties.}
Nagoya Mathematical Journal \textbf{102} (1986), 79--90.

\bibitem{Sul}
D. Sullivan.
\textit{Infinitesimal computations in topology.}
Publications Mathématiques de l’Institut des Hautes Études Scientifiques \textbf{47} (1977), 269--331.
\end{thebibliography}

\documentclass[11pt]{amsart}
\usepackage[T1]{fontenc}
\usepackage{amsmath,amssymb,amsthm}
\usepackage{mathtools}
\usepackage[hypertexnames=false]{hyperref}%ないとだめ
\usepackage{cleveref}
\usepackage{autonum}
\usepackage{booktabs}
\usepackage{longtable}
\usepackage{float}
\newtheorem{thm}{Theorem}[section]
\newtheorem{lem}{Lemma}[section]
\newtheorem{prop}{Proposition}[section]
\newtheorem{cor}{Corollary}[section]
\newtheorem{eg}{Example}[section]
\newtheorem{definition}{Definition}[section]
\newtheorem{q}{Question}[section]
\newtheorem{rem}{Remark}[section]

\newtheorem{claim}{Claim}[section]
\newtheorem{prob}{Problem}[section]

\crefname{thm}{Theorem}{Theorem}
\crefname{lem}{Lemma}{Lemma}
\crefname{prop}{Proposition}{Proposition}
\crefname{cor}{Corollary}{Corollary}
\crefname{eg}{Example}{Example}
\crefname{definition}{Definition}{Definition}
\crefname{q}{Question}{Question}
\crefname{rem}{Remark}{Remark}
\crefname{conj}{Conjecture}{Conjecture}
\crefname{claim}{Claim}{Claim}
\crefname{prob}{Problem}{Problem}

\def\fr#1{\mathfrak{#1}}
\def\bb#1{\mathbb{#1}}

\def\ali#1{\begin{align}#1\end{align}}
\def\dis#1{\displaystyle{#1}}
\def\w{\wedge}
\def\bw{\bigwedge}
\def\o{\oplus}
\def\bo{\bigoplus}
\def\ang#1{\langle #1\rangle}
\def\dim#1{{\rm dim\ }#1}
\def\0{\{ 0\}}
\def\Ker#1{{\rm Ker}\ #1}
\def\Im#1{{\rm Im}\ #1}

\title[The nilpotent funfdamental groups]{On the torsion-free nilpotent fundamental groups of smooth quasi-projective varieties of rank up to seven}
\author{Taito Shimoji}
\date{}
\address{Department of Mathematics, Graduate School of Science, The University of Osaka, Osaka, Japan}
\email{u215629i@ecs.osaka-u.ac.jp}
\begin{document}
\maketitle
\begin{abstract}
Let $X$ be a smooth quasi-projective variety. Assume that the (topological) fundamental group $\pi_1(X, x)$ is torsion-free nilpotent. We show that if the first Betti number $b_1(X) \le 3$, then $\pi_1(X, x)$ is isomorphic to either $\mathbb{Z}^n$ for $n = 1, 2, 3$, a lattice in the Heisenberg group $H_3(\mathbb{R})$ or $\mathbb{R} \times H_3(\mathbb{R})$. Moreover, we prove that $\pi_1(X, x)$ is abelian or $2$-step nilpotent if its rank is less than or equal to seven. More precisely, we determine the real nilpotent Lie groups in which torsion-free nilpotent fundamental groups can be embedded as lattices for ranks up to six and seven, respectively. Our main results are a partial positive answer to a question on nilpotent (quasi-)K\"ahler groups posed by Aguilar and Campana.
\end{abstract}
\section{Introduction}
\subsection{Aguilar and Campana's question for the nilpotent fundamental groups of smooth quasi-projective varieties}

The following question is given by Aguilar and Campana in \cite{RodCam}.
\begin{q}
Let $G$ be a torsion-free nilpotent Kähler group. Can its
 nilpotency class $\nu (G)$ be at least $3$? Can $\nu (G)$ be arbitrarily large?
 Same questions for the quasi-Kähler groups.
\end{q}
This question is mainly motivated by the previous works of Campana and others (for example, see \cite{Cam},\cite{Cam2},\cite{AraNori} and \cite{CadDengYam},etc). We focus on the case that algebraic varieties are smooth quasi-projective. Thus we consider the following problem. 
\begin{prob}\label{Campana}
Let $X$ be a smooth quasi-projective variety over $\bb{C}$. Assume that the (topological) fundamental group $\pi_1(X,x)$ is torsion-free nilpotent. Is $\pi_1(X,x)$ abelian or $2$-step nilpotent?
\end{prob}
In fact, all the known examples of the nilpotent fundamental groups of smooth quasi-projective varieties are either abelian or $2$-step nilpotent (see \cite{Pierre}, \cite{Cam},\cite{JohVen} and \cite{CadDengYam}). 
 
\subsection{Main theorems}
Let $X$ be a smooth quasi-projective variety over $\bb{C}$. The first betti number $b_1(X)$ is given by 
\ali{b_1(X):=\dim H^1(X,\bb{R})}
where $H^1(X,\bb{R})$ is the first real cohomology of $X$.
The first main theorem is as follows:
\begin{thm}[\protect\cref{b1<=3}]
Let $X$ be a smooth quasi-projective variety over $\bb{C}$. If $b_1(X)\leq 3$ and $\pi_1(X,x)$ is torsion-free nilpotent, then $\pi_1(X,x)$ is abelian or $2$-step nilpotent. More precisely, if $b_1(X)\leq 3$ and $\pi_1(X,x)$ is torsion-free nilpotent, then $\pi_1(X,x)$ is isomorphic to $\bb{Z}$, $\bb{Z}^2$, $\bb{Z}^3$, a lattice in $H_3(\bb{R})$ or $H_3(\bb{R})\times\bb{R}$ where $H_3(\bb{R})$ is $3$-dimensional Heisenberg group.
\end{thm}
Let $\Gamma$ be a torsion-free $s$-step nilpotent group. A \textit{rank} $rk(\Gamma )$ of $\Gamma$ is defined as 
\ali{
rk(\Gamma ):=\sum_{i=1}^{s}rk(C^i\Gamma /C^{i+1}\Gamma )
}
where $C^i\Gamma$ is the lower central series and $C^{s+1}\Gamma=\{ e\}$. Note that \ali{
b_1(X)\leq rk(\pi_1(X,x)).
}
Next we study the case of higher betti numbers. By restricting the rank of $\pi_1(X,x)$, we obtain the second main theorem concerning \cref{Campana}.   
\begin{thm}[\cref{rk<=7}]\label{main3}
Let $X$ be a smooth quasi-projective variety. Assume that $\pi_1(X,x)$ is torsion-free nilpotent. If $rk(\pi_1(X,x))\leq 7$, then $\pi_1(X,x)$ is abelian or $2$-step nilpotent.
\end{thm}%rank6までは分類できるというcorを述べる。５次元ハイゼンベルグも忘れずに
The proof of the above theorem follows from the condition of the bigraded Lie algebras such that it is shown by Morgan (\cite[Theorem {\bf (9.4)}]{Morgan}). Using the condition of bigraded Lie algebras, we determine the simply-connected real nilpotent Lie groups in which the torsion-free nilpotent fundamental groups can be embedded as lattices. The followings \cref{prop1} and \cref{prop2} are the cases of $rk(\pi_1(X,x))\leq 6$ and $rk(\pi_1(X,x))=7$ respectively.
\begin{thm}[\cref{rk<=6}]\label{prop1}
Let $X$ be a smooth quasi-projective variety. If $\pi_1(X,x)$ is torsion-free nilpotent and $rk(\pi_1(X,x))\leq 6$, then $\pi_1(X,x)$ is a lattice in one of the following simply-connected nilpotent Lie groups:%すべてのlatticeではない
\ali{
\bullet\ &\bb{R}&\text{ if }b_1(X)=1,\\
\bullet\ &\bb{R}^2\text{ or }H_3(\bb{R})&\text{ if }b_1(X)=2,\\
\bullet\ &\bb{R}^3\text{ or }\bb{R}\times H_3(\bb{R})&\text{ if }b_1(X)=3,\\
\bullet\ &\bb{R}^4, \bb{R}^2\times H_3(\bb{R}),H_3(\bb{R})\times H_3(\bb{R})\text{ or }H_5(\bb{R})&\text{ if }b_1(X)=4,\\
\bullet\ &\bb{R}^5, \bb{R}^3\times H_3(\bb{R})\text{ or }\bb{R}\times H_5(\bb{R})&\text{ if }b_1(X)=5,\\
\bullet\ &\bb{R}^6&\text{ if }b_1(X)=6 
}  
where $H_{2k+1}(\bb{R})$ is $(2k+1)$-dimensional Heisenberg group.
\end{thm}
\begin{thm}[\cref{rk=7}]\label{prop2}
Let $X$ be a smooth quasi-projective variety. If $\pi_1(X,x)$ is torsion-free nilpotent and $rk(\pi_1(X,x))=7$, then $\pi_1(X,x)$ is isomorphic to a lattice in a simply-connected real  nilpotent Lie group such that the complexification $\fr{n}_{\bb{C}}$ of the Lie algebra $\fr{n}$ of $N$ is isomorphic to one of the following Lie algebras: 
\ali{
\bullet\ &\bb{C}^7&\text{ if }b_1(X)=7,\\
\bullet\ &\bb{C}^4\o\fr{n}_3(\bb{C}), \bb{C}^2\o L_{5,4} \text{ or }\fr{n}_7^{152}&\text{ if }b_1(X)=6,\\
\bullet\ &\bb{C}\o\fr{n}_3(\bb{C})\o\fr{n}_3(\bb{C})\text{ or }(\text{other $2$-step nilpotent such that $\dim H^1(\fr{n})=5$})&\text{ if }b_1(X)=5,\\
\bullet\ &\fr{n}_7^{142}\text{ or }\fr{n}_7^{143}&\text{ if }b_1(X)=4
}
where $L_{5,4}$ is the complexification of the Lie algebra of $H_5(\bb{R})$, $\fr{n}_7^{152}$ is the complexification of the Lie algebra of $H_7(\bb{R})$, $\fr{n}_7^{142}$ and $\fr{n}_7^{143}$ are given by
\ali{
\fr{n}_7^{142}=\ang{X_1,\dots,X_7}: [X_1,X_i]=X_{i-1}\text{ for i=3,5,7 },[X_3,X_5]=X_4,[X_5,X_7]=X_2,
}
and
\ali{
\fr{n}_7^{143}=\ang{X_1,\dots,X_7}: [X_1,X_i]=X_{i-1}\text{ for i=3,5,7 },[X_3,X_5]=X_6,[X_5,X_7]=X_2
.}
\end{thm}
Note that the Lie algebras $\fr{n}_7^{142}$ and $\fr{n}_7^{143}$ are not isomorphic to $(2k+1)$-dimensional Heisenberg Lie algebra for any $k$. 
\begin{rem}[\cref{8dim}]
If $rk(\pi_1(X,x))=8$ and $\pi_1(X,x)$ is torsion-free nilpotent, then we cannot conclude whether \cref{Campana} holds by using the method of the proof of main theorems. In particular, there exists an example of a $3$-step nilpotent group which may be realized as the fundamental group of a smooth quasi-projective variety. The example is a lattice in the following real Lie group:
\ali{
N(\bb{R}):=\left \{\begin{pmatrix}
1&0&y_2&y_1&-z_2&-z_1&0&\frac{1}{2}(y_1+y_2)&3b\\
0&1&x_2&x_1&0&\frac{1}{2}(x_1+x_2)&-z_2&-z_1&3a\\
0&0&1&0&x_1&x_2&-y_1&-y_2&2z_2\\
0&0&0&1&x_2&x_1&-y_2&-y_1&2z_1\\
0&0&0&0&1&0&0&0&y_2\\
0&0&0&0&0&1&0&0&y_1\\
0&0&0&0&0&0&1&0&x_2\\
0&0&0&0&0&0&0&1&x_1\\
0&0&0&0&0&0&0&0&1\\
\end{pmatrix}
;
x_1,x_2,y_1,y_2,z_1,z_2,a,b\in\bb{R}\right \}
.
}
Moreover, the Lie algebra of $N(\bb{R})$ is given by
\ali{
\fr{g}:=\ang{X_1,X_2,Y_1,Y_2,Z_1,Z_2,A,B}:&[X_1,Y_1]=Z_1=[X_2,Y_2], [X_2,Y_1]=Z_2=[X_1,Y_2],\\
&[X_1,Z_1]=A=[X_2,Z_2],[Y_1,Z_1]=B=[Y_2,Z_2]
.}
\end{rem}
\subsection{Acknowledgements}
The author thanks his supervisors, Professor Hisashi Kasuya, Nagoya University, for his enormous support and pointing out how to state the main theorems. He is also grateful to Professor Katsutoshi Yamanoi,  The University of Osaka, for his valuable comments regarding the example in \cref{8dim}.
The author also benefited greatly from Rogov’s paper \cite{Rogov}, which was particularly helpful for \cref{Campana}. This work was supported by JST SPRING, Grant Number JPMJSP2138.
\section{Mixed Hodge theory on nilpotent fundamental groups of smooth quasi-projective varieties} 
\subsection{Mixed Hodge structures}(\cite{Del},\cite{Morgan},\cite{PeterSteen})
An  $\bb{R}$-\textit{mixed Hodge structure} $(V,W,F)$ is defined by the following data:
\begin{itemize}
\item $V$ is an $\bb{R}$-vector space.
\item $W=\{W_k(V)\}_{k\in\bb{Z}}$ is a finite increasing filtration on $V$.
\item $F=\{F^p(V_{\bb{C}})\}_{p\in\bb{Z}}$ is a finite decreasing filtration on the complexification $V_{\bb{C}}:=V\otimes\bb{C}$ of $V$ and defines a Hodge structure on 
\ali{
Gr_k^W(V_\bb{C}):=\frac{W_k(V)\otimes\bb{C}}{W_{k-1}(V)\otimes\bb{C}}
}
of weight $k$ for all $k\in\bb{Z}$. The decreasing filtration is given by
\ali{
F^p(Gr_k^W(V_\bb{C}))=\frac{F^p(V_{\bb{C}})\cap (W_k\otimes \bb{C})}{W_{k-1}\otimes\bb{C}}.
}
{\it i.e.} there exists a bigrading $Gr_k^W(V_{\bb{C}})=\bo_{p+q=k}\mathcal{H}_{p,q}$ such that
\ali{
\overline{\mathcal{H}_{p,q}}=\mathcal{H}_{q,p}\ \text{ for all }p,q\in\bb{Z}\text{ with }p+q=k.
}
\end{itemize}
Let $(V,W,F)$ be a mixed Hodge structure. Equivalently, we have the bigrading $V_{\bb{C}}=\bo_{p,q}V_{p,q}$ such that
\ali{
\overline{V_{p,q}}\equiv V_{q,p}\mod \bo_{s+t<p+q}V_{s,t}.
}
Here, $V_{p,q}$ is given by
\ali{
V_{p,q}:=F^p(V_{\bb{C}})\cap W_{p+q}(V_{\bb{C}})\cap \left (\overline{F^q(V_{\bb{C}})}\cap W_{p+q}(V_{\bb{C}})+\sum_{i\geq 2}\overline{F^{q-i+1}(V_{\bb{C}})}\cap W_{p+q-i}(V_{\bb{C}})\right)}
and $W_k(V_{\bb{C}}):=W_k\otimes\bb{C}$. Moreover, the bigrading satisfies
\ali{
W_k(V_{\bb{C}})&=\bo_{p+q\leq k}V_{p,q}
}
and
\ali{
F^p(V_{\bb{C}})&=\bo_{s\geq p,t\in\bb{Z}}V_{s,t}
.}
\subsection{Mixed Hodge structures on Lie algebras}
\begin{definition}
Let $\fr{g}$ be an $\bb{R}$-Lie algebra. The bigrading of a mixed Hodge structure on $\fr{g}$ is a bigrading $\fr{g}_{\bb{C}}=\bo_{p,q\in \bb{Z}}\fr{g}_{p,q}$ satisfying
\ali{
\overline{\fr{g}_{p,q}}\equiv \fr{g}_{q,p}\mod \bo_{k+l<p+q}\fr{g}_{k,l}
}and
\ali{
[\fr{g}_{p,q},\fr{g}_{s,t}]\subset \fr{g}_{p+s,q+t}
}
for all $p,q,s,t\in\bb{Z}$. In particular, the filtrations $F$ on $V_{\bb{C}}$ and $W$ on $V_{\bb{C}}$ are given by
\ali{
W_k(\fr{g}_{\bb{C}})&=\bo_{p+q\leq k}\fr{g}_{p,q}
}
and
\ali{
F^p(\fr{g}_{\bb{C}})&=\bo_{s\geq p,t\in\bb{Z}}\fr{g}_{s,t}.
}
\end{definition}
Let $\fr{g}_{\bb{C}}=\bo_{p,q\in \bb{Z}}\fr{g}_{p,q}$ be the bigrading of a mixed Hodge structure on a Lie algebra $\fr{g}$. By the compatibility with the Lie bracket of $\fr{g}$, we have the bigrading of a mixed Hodge structure on the Lie algebra cohmology $H^*(\fr{g}_{\bb{C}})=\bo_{p,q\in\bb{Z}}H_{p,q}^*$. More precisely, we define a subspace $C_{p,q}^j\subset \bw^j\fr{g}_{\bb{C}}^*$ to be 
\ali{
C_{p,q}^j:=\bo_{\substack{p=-(p_1+\cdots +p_j)\\ q=-(q_1+\cdots +q_j)}}\fr{g}_{p_1,q_1}^*\w\cdots\w\fr{g}_{p_j,q_j}^*
}
where $\fr{g}_{p_1,q_1}^*\w\cdots\w\fr{g}_{p_j,q_j}^*$ is a subvector space generated by the set 
\ali{
\{ f_1\w\cdots\w f_j\mid f_i\in \fr{g}_{p_i,q_i}^*\}.
} 
Then we can show the following equalities: 
\ali{
\bw ^j\fr{g}_{\bb{C}}^*&=\bo_{p,q\in\bb{Z}}C_{p,q}^j,\\
d(C_{p,q}^j)&\subset C_{p,q}^{j+1}
}
and
\ali{
H^j(\fr{g}_{\bb{C}})&=\bo_{s,t\in\bb{Z}}H_{s,t}^j
}
where $d$ is the differential of $\fr{g}_{\bb{C}}$ and 
\ali{
H_{s,t}^j:=\dis{\frac{\Ker{(d:C_{s,t}^j\rightarrow C_{s,t}^{j+1})}}{\Im{(d:C_{s,t}^{j-1}\rightarrow C_{s,t}^{j})}}}.
}
Given a basis for each homogeneous component of a bigraded vector space, for instance, we denote the degree $(p,q)$ component, whose basis consists of $X_1\dots X_n$ by $\ang{X_1,\dots ,X_n}_{p,q}.$ 
\begin{eg}\label{Heisenberg}
Let $\fr{n}_3(\bb{R})$ the $3$-dimensional Heisenberg Lie algebra. We can take a basis of the complexification $\fr{n}_3(\bb{R})\otimes\bb{C}=\fr{n}_3(\bb{C})=\ang{X_1,X_2,X_3}$ such that $[X_1,X_2]=X_3,[X_1,X_3]=0=[X_2,X_3]$ and $X_2=\overline{X_1}$. We define the bigrading on $\fr{n}_3(\bb{C})$ by
\ali{
\fr{n}_3(\bb{C})=\ang{X_1}_{-1,0}\o\ang{X_2}_{0,-1}\o\ang{X_3}_{-1,-1}
.}
The bigrading is a mixed Hodge structure on $\fr{n}_3(\bb{R})$. The induced bigradings on the cohomologies are given by
\ali{
H^1(\fr{n}_3(\bb{C}))&=H_{1,0}^1\o H_{0,1}^1=\ang{[x_1]}_{1,0}\o\ang{[x_2]}_{0,1},\\
H^2(\fr{n}_3(\bb{C}))&=H_{2,1}^2\o H_{1,2}^2=\ang{[x_1\w x_3]}_{2,1}\o\ang{[x_2\w x_3]}_{1,2},\\
H^3(\fr{n}_3(\bb{C}))&=H_{2,2}^3=\ang{[x_1\w x_2\w x_3]}_{2,2}
}
where $x_1,x_2$ and $x_3$ are the dual basis of $X_1,X_2$ and $X_3$.
\end{eg}
\subsection{Bigradings of mixed Hodge structures and a restriction on the nilpotent fundamental groups of smooth quasi-projective varieties}
Let $X$ be a smooth quasi-projective variety. Assume that $\Gamma =\pi_1(X,x)$ is torsion-free nilpotent. By \cite{Malcev}, there uniquely exists a simply-connected nilpotent Lie group $N$ such that $\Gamma$ is a lattice (discrete and cocompact subgroup) in $N$. In \cite{Morgan}, Morgan showed the necessary condition of the Lie algebra $\fr{n}$ of $N$ in terms of the bigrading of a mixed Hodge structure on $\fr{n}$. More precisely, Morgan showed the following theorem.
\begin{thm}\cite[Theorem {\bf (9.4)}]{Morgan}\label{W}
Let $X$ be a smooth quasi-projective variety. Let $N$ be a simply-connected real nilpotent Lie group. Assume that $\pi_1(X,x)$ is a lattice in $N$. For the Lie algebra $\fr{n}$ of $N$, there exists the bigrading $\fr{n}_{\bb{C}}=\bo_{p,q\leq 0,p+q\leq -1}\fr{n}_{p,q}$ of a mixed Hodge structure such that for the induced bigrading $H^*(\fr{g}_{\bb{C}})=\bo_{p,q\geq 0,p+q\geq 1}H_{p,q}^*$, we have 
\ali{
{\bf (W)}: H^1(\fr{n}_{\bb{C}})&=H_{1,0}^1\o H_{0,1}^1\o H_{1,1}^1,\\
H^2(\fr{n}_{\bb{C}})&=H_{2,0}^2\o H_{1,1}^2\o H_{0,2}^2\o H_{2,1}^2\o H_{1,2}^2\o H_{2,2}^2.
}
\end{thm}
The condition {\bf (W)} is derived from the bigradings of canonical mixed Hodge structures on the cohomologies of complex algebraic varieties (see \cite{Del}). Consequently, we have
\begin{cor}
Let $\Gamma$ be a lattice in a simply-connected nilpotent Lie group $N$. If the complexification of the Lie algebra $\fr{n}_{\bb{C}}$ of N does not have the bigrading of any mixed Hodge structure satisfying the condition {\bf (W)}, then $\Gamma$ is not isomorphic to the fundamental group of any smooth quasi-projective variety.
\end{cor}
\begin{eg}\label{abelians}
Let $\bb{C}$ be the $1$-dimensional abelian Lie algebra. Take a non-zero element $X$. Then the bigrading $\bb{C}=\ang{X}_{-1,-1}$ satisfies the conditions {\bf (W)}. In the case of the $2$-dimensional abelian Lie algebra $\bb{C}^2$, we may set $\bb{C}^2=\ang{X}_{-1,0}\o\ang{\overline{X}}_{0,-1}$ where $X$ and $\overline{X}$ form a basis of $\bb{C}^2.$  
\end{eg}
\begin{prop}\label{trivialextension}
Let $\fr{n}$ be a $\bb{C}$-Lie algebra which admits the bigrading of a mixed Hodge structure in \cref{W}. Then any trivial extension $\fr{n}\o\bb{C}^m$ also admits such a bigrading.
\end{prop}
\begin{proof}
Let $\fr{n}=\bo_{p,q\leq 0, p+q\leq -1}\fr{n}_{p,q}$ be the bigrading of a mixed Hodge structure satisfying the condition {\bf (W)}, and the induced grading $H^*(\fr{n})=\bo_{p,q\geq 0, p+q\geq 1}H_{p,q}^*$. We define $({\fr{n}\o\bb{C}^m})_{-1,-1}:=\fr{n}_{-1,-1}\o\bb{C}^m$ and $({\fr{n}\o\bb{C}^m})_{p,q}:=\fr{n}_{p,q}$ for other $i$. Since the differentiation of the dual basis for any basis of $\bb{C}^m$ are all zero, we have 
\ali{d(x\w y)=dx\w y}
for all $x\in \bw^{\bullet}\fr{n}^*$ and $y\in \bw^{\bullet}(\bb{C}^m)^*.$ Hence we have $dx=0$ if and only if $d(x\w y)=0$ for any $x\in \bw^{\bullet}\fr{n}^*$ and $y\in \bw^{\bullet}(\bb{C}^m)^*.$ Thus we see that 
\ali{
H^1(\fr{n}\o\bb{C}^m)=H_{1,0}^1\o H_{0,1}^1\o {H_{1,1}^1}'
}
\ali{H^2(\fr{n}_{\bb{C}})=H_{2.0}^2\o H_{1,1}^2\o H_{0,2}^2\o {H_{1,2}^2}'\o {H_{2,1}^2}'\o{ H_{2,2}^2}'
}
where ${H_{1,1}^1}'=H_{1,1}^2\o (\bb{C}^m)^*$, ${H_{1,2}^2}'=H_{1,2}^2\o\ang{[x\w y]\mid x\in (\fr{n}_{0,-1})^*,y\in (\bb{C}^m)^*},$ ${H_{2,1}^2}'=H_{2,1}^2\o\ang{[x\w y]\mid x\in (\fr{n}_{-1,0})^*,y\in (\bb{C}^m)^*}$ and ${H_{2,2}^2}'=H_{2,2}^2\o\ang{[y_i\w y_j]\mid y_i,y_j\in (\bb{C}^m)^*, 1\leq i<j\leq m}.$ Thus the grading satisfies the condition {\bf (W)}.   
\end{proof}
\section{Main theorem:the case of $b_1(X)\leq 3$}\label{9}
We denote by $\mathcal{M}$ the class of nilpotent Lie algebras over $\bb{R}$ such that its complexification admit bigradings of mixed Hodge structures in \cref{W}. 

Let $X$ be a smooth quasi-projective variety. We denote the first betti number of $X$ by $b_1(X)=\dim H^1(X,\bb{R})$. In this section, we show the following theorem:
\begin{thm}\label{b1<=3}
Let $X$ be a smooth quasi-projective variety. If $b_1(X)\leq 3$ and $\pi_1(X,x)$ is torsion-free nilpotent, then $\pi_1(X,x)$ is isomorphic to $\bb{Z}$,$\bb{Z}^2$,$\bb{Z}^3$, a lattice in $H_3(\bb{R})$ or $H_3(\bb{R})\times\bb{R}$ where $H_3(\bb{R})$ is $3$-dimensional Heisenberg group. In particular, if $b_1(X)\leq 3$, then $\pi_1(X,x)$ is abelian or $2$-step.
\end{thm}
Since $\Gamma =\pi_1(X,x)$ is torsion-free nilpotent, there uniquely exists a simply-connected nilpotent Lie group $N$ such that $\Gamma$ is a lattice in $N$ (\cite{Malcev},\cite{Raghu}). In this case, by \cite{Morgan}, The complex $(\bw^{\bullet}\fr{n}^*,d)$ of the Lie algebra $\fr{n}$ of $N$ is a $1$-minimal model of $X$ in the sense of Sullivan(\cite{Sul}). Thus $b_1(X)=\dim H^1(X,\bb{R})=\dim H^1(\fr{n})$. Thus it suffices to study the Lie algebra.
\subsection{The Case of Two Generators ($b_1(X)=2$)}
\begin{prop}\label{twogen}
Let $X$ and $\overline{X}$ be elements of $\fr{g}_{\bb{C}}$ such that their dual elements are  $x$ and $\overline{x}$ respectively. If $\fr{g}\in\mathcal{M}$ and $\dim H^1(\fr{g}_{\bb{C}}) = 2$, then $\fr{g}_{\bb{C}}$ is either isomorphic to $\bb{C}^2$ or the Heisenberg Lie algebra $\fr{n}_3(\bb{C})$. 
\end{prop}
\begin{proof}
Let $\fr{g}\in\mathcal{M}$ and $\fr{g}_{\bb{C}}=\bo_{p,q\leq 0, p+q\leq -1}\fr{g}_{p,q}$. We assume that $\dim H^1(\fr{g}_{\bb{C}})=2$ and $\dim \fr{g}_{\bb{C}}>3$. Since $\fr{g}\in\mathcal{M}$, the induced gradings is either
\ali{
H^1(\fr{g})=H_{1,0}^1\o H_{0,1}^1=\ang{[x]}_{1,0}\o\ang{[\overline{x}]}_{0,1}
}
or
\ali{
H^1(\fr{g})=H_{1,1}^1=\ang{[x],[\overline{x}]}_{1,1}.
}
Assume that $\fr{g}_{-1,-1}=\ang{X,\overline{X}}_{-1,-1} (\text{resp. }\fr{g}_{-1,0}=\ang{X}_{-1,0}\text{ and }\fr{g}_{0,-1}=\ang{\overline{X}}_{0,-1})$ For all integer $p\geq 2$, there exists an element $z_p\in \fr{g}_{-p,-p}($ resp. $\fr{g}_{-1,-p}$ or $\fr{g}_{-p,-1})$ such that
\ali{
dz_p=x\w z_{p-1}
}
or
\ali{
dz_p=\overline{x}\w z_{p-1}.
}
We prove by induction for $p\geq 2$. Since $\dim\fr{g}>3$, $Z_2=[X,\overline{X}]\neq 0$. Let $x,\overline{x}$ and $z_2$ be the dual element of $X,\overline{x}$ and $Z_2$. Let $z_1:=-\overline{x}$. Thus we have
\ali{
dz_2=-x\w \overline{x}=x\w z_1
.}
Assume that there exists $z_p\in \fr{g}_{-p,-p}^*$ such that $dz_p=x\w z_{p-1}$. Since $d(x\w z_p)=0, x\w z_p\in C_{p+1,p+1}^2, p+1\geq 3\text{ and }\fr{g}\in\mathcal{M}$,
there exists $z_{p+1}\in \fr{g}_{-(p+1),-(p+1)}^*$ such that 
\ali{
dz_{p+1}=x\w z_p
.}
Therefore the lemma holds. When $\fr{g}_{-1,0}=\ang{X}_{-1,0}\text{ and }\fr{g}_{0,-1}=\ang{\overline{X}}_{0,-1}$, we can show by the same argument.
Thus $\fr{g}\notin\mathcal{M}$ if $\dim\fr{g}>3$. Therefore $\dim\fr{g}\leq 3$ and $\fr{g}_{\bb{C}}$ is either isomorphic to $\bb{C}^2$ or $\fr{n}_3(\bb{C})$.
\end{proof}
Since $\dim H^1(\fr{n})=2$ if $\fr{n}$ is filiform, we obtain the following.
\begin{cor}[\cite{grading}]
Let $\Gamma$ be a lattice in a simply-connected nilpotent Lie group $N$. and $\fr{n}$ the Lie algebra of $N$. Assume that $\fr{g}$ is {\it filiform}. $\Gamma$ is isomorphic to the fundamental group of a smooth quasi-projective variety if and only if $\dim \fr{n}=2$ or $3$.
\end{cor}
\subsection{The Case of Three Generators ($b_1(X)=3$)}
\begin{prop}\label{threegen}
Let $\fr{g}\in\mathcal{M}$ and $\fr{g}_{\bb{C}}=\bo_{p,q\leq 0, p+q\leq -1}\fr{g}_{p,q}\in \mathcal{M}$. If $\dim H^1(\fr{g}_{\bb{C}})=3$, then $\fr{g}_{\bb{C}}$ is isomorphic either to $\bb{C}^3$ or $\fr{n}(\bb{C})\o \bb{C}$. 
\end{prop}
\begin{proof}
Let $\fr{g}\in\mathcal{M}$ and $\fr{g}_{\bb{C}}=\bo_{p,q\leq 0, p+q\leq -1}\fr{g}_{p,q}\in \mathcal{M}$. We assume that $\dim H^1(\fr{g}_{\bb{C}})=3$ and $\dim \fr{g}_{\bb{C}}>4$. Since $\fr{g}\in\mathcal{M}$, the induced gradings is either
\ali{
H^1(\fr{g})=H_{1,0}^1\o H_{0,1}^1\o H_{1,1}^1=\ang{[x]}_{1,0}\o\ang{[\overline{x}]}_{0,1}\o\ang{[y]}_{1,1}
}
or
\ali{
H^1(\fr{g})=H_{1,1}^1=\ang{[x],[y],[z]}_{1,1}\text{ such that }\overline{H_{1,1}^1}=H_{1,1}^1
.}
Let $\fr{g}_{-1,0}=\ang{X}_{-1,0},\fr{g}_{0,-1}=\ang{\overline{X}}_{0,-1}\text{ and }\fr{g}_{-1,-1}=\ang{Y}_{-1,-1}$ such that $x,\overline{x}$ and $y$ are the dual elements. We first assume that $Z_1:=[X,\overline{X}]\neq 0$. Since $\dim\fr{g}>4$, we have 
\ali{
Z_2:=[X,Z_1]\neq 0\ ([\overline{X},Z_1]\neq 0)
}
or
\ali{
Z_3:=[X,Y]\neq 0\ ([\overline{X},Y]\neq 0).
}
If $Z_2\neq 0$, then we can show that $\fr{g}_{-p,-1}\neq \0$ for all $p\geq 2$, and hence $\dim\fr{g}=\infty$. Thus $Z_2=0$ and $Z_3\neq 0$. As the case $Z_2\neq 0$, we can show that $\fr{g}_{-p,-1}\neq \0$ for all $p\geq 2$, and hence $\fr{g}\notin\mathcal{M}$. We assume that $\fr{g}_{-1,-1}=\ang{X,Y,Z}_{-1,-1}$ such that $\overline{\fr{g}_{-1,-1}}=\fr{g}_{-1,-1}$. Since $\dim\fr{g}>4$, we have $\fr{g}_{-2,-2}\neq \0$. We can easily show that $\fr{g}_{-p,-p}\neq \0$ for all $p\geq 3$ if $\dim\fr{g}_{-2,-2}=1$ or $2$. We assume that $\dim\fr{g}_{-2,-2}=1$. We may assume that $Z_1:=[X,Y]$ is a basis of $\fr{g}_{-2,-2}$. Since $\dim C_{3,3}^3=\dim\ang{x\w y\w z}=1$, we have 
\ali{
\dim\Ker (d: C_{3,3}^2\rightarrow C_{3,3}^3)=2,
}
and hence $\dim\fr{g}_{-3,-3}=2$. In addition, we can show that $\fr{g}_{-p,-p}\neq \0$ for all $p\geq 3$ by induction. Thus $\fr{g}\notin\mathcal{M}$ and $\dim\fr{g}\leq 4$. Therefore $\fr{g}_{\bb{C}}$ is either isomorphic to $\bb{C}^3$ or $\fr{n}_3(\bb{C})\o\bb{C}$. 
\end{proof}
\section{Main theorem:The case of $rk(\pi_1(X,x))\leq 7$}
In this section, we show the following theorem.
\begin{thm}\label{rk<=7}
Let $X$ be a smooth quasi-projective variety such that $\pi_1(X,x)$ is torsion-free nilpotent and $rk(\pi_1(X,x))\leq 7$. Then $\pi_1(X,x)$ is abelian or $2$-step. 
\end{thm}
\subsection{The proof of \cref{rk<=7}}
We first show the following lemma. The following is the case of $b_1(X)=4$ and $3$-step.
\begin{lem}\label{geq8}
Let $N$ be a a simply-connected nilpotent Lie group such that $\pi_1(X,x)$ is isomorphic to a lattice in $N$. Let $\fr{g}_{\bb{C}}$ be the complexification of the Lie algebra of $N$. If $b_1(X)=4$ and $\fr{g}_{\bb{C}}$ is $3$-step nilpotent, then $\dim\fr{g}_{\bb{C}}\geq 8$.
\end{lem}
\begin{proof}
We define $(x,y,z):=(\dim H_{1,0}^1,\dim H_{0,1}^1,\dim H_{1,1}^1)$. Assume that $\dim \fr{g}\leq 7$. 

\underline{$(1):$ $(x,y,z)=(2,2,0)$}

Let $\fr{g}_{-1,0}=\ang{X_1,X_2}\text{ and }\fr{g}_{0,-1}=\ang{\overline{X_1},\overline{X_2}}$.
Then we can show that $[X_1,X_2]=0=[\overline{X_1},\overline{X_2}]$ by the same argument as in \cref{twogen}. Assume that $\dim \fr{g}_{-1,-1}=2$. Let $\fr{g}_{-1,-1}=\ang{Z_1,\overline{Z_1}}\subset C^1\fr{g}_{\bb{C}}$. Since $\fr{g}_{\bb{C}}$ is $3$-step and $[Z_1,Z_2]\in C^3\fr{g}=\0$, we have $[X_i,Z_j]\neq 0$ for some $i,j$. Thus $\dim \fr{g}_{-2,-1}=\dim\fr{g}_{-1,-2}\neq 0$. This shows that $\dim\fr{g}\geq 8$. Thus $\dim\fr{g}_{-1,-1}=1$. Let $\fr{g}_{-1,-1}=\ang{Z_1}\subset C^1\fr{g}_{\bb{C}}$ such that $\overline{\fr{g}_{-1,-1}}=\fr{g}_{-1,-1}$. Since $\fr{g}_{\bb{C}}$ is $3$-step, we have $[X_i,Z_1],[\overline{X_j},Z_1]\neq 0$ for some $i,j$. Furthermore $\dim\fr{g}\leq 7$, we have $\dim\fr{g}_{-2,-1}=\dim\fr{g}_{-1,-2}=1$. Let $\fr{g}_{-2,-1}=\ang{Z_2}$ and $\fr{g}_{-1,-2}=\ang{Z_3}$. Let $x_1,x_2,\overline{x_1},\overline{x_2},z_1,z_2$ and $z_3$ be the dual basis of $X_1,\dots ,Z_3$. We write
\ali{
dz_2=a_1x_1\w z_1+a_2x_2\w z_1=(a_1x_1+a_2x_2)\w z_1\text{ and }(a_1,a_2)\neq (0,0). 
}
We set $x:=a_1x_1+a_2x_2\neq 0$. Then $d(x\w z_2)=0$ and $x\w z_2\in C_{3,1}^2\cap\Ker d$. Since $\fr{g}\in\mathcal{M}$, there exists $0\neq f\in \fr{g}_{-3,-1}^*$ such that $df=x\w z_2$. This contradicts that $\dim\fr{g}\leq 7$. Thus we have $(x,y,z)=(1,1,2)$ or $(0,0,4)$. 

\underline{$(2):$ $(x,y,z)=(1,1,2)$} 

We can show that the existence of non-zero elements in $H_{3,1}^2$ and $H_{1,3}^2$ by the same argument as in the case that $(x,y,z)=(2,2,0)$.

\underline{$(3):$ $(x,y,z)=(0,0,4)$}

We can show that the existence of non-zero elements in $H_{3,3}^2$ when $\dim\fr{g}_{-2,-2}=1$ by the same argument as in the case that $(x,y,z)=(2,2,0)$. Thus $\dim\fr{g}_{-2,-2}=2$. Let $\fr{g}_{-2,-2}=\ang{Z_1,Z_2}$. Since $C_{3,3}^2=\ang{x_i\w z_j\mid 1\leq i\leq 4,j=1,2}$ and $C_{3,3}^3=\ang{x_i\w x_j\w x_k\mid 1\leq i<j<k\leq 4}$, we have
\ali{
\dim\Ker (d:C_{3,3}^2\rightarrow C_{3,3}^3)&=\dim C_{3,3}^2-\dim \Im (d:C_{3,3}^2\rightarrow C_{3,3}^3)\\
&\geq \dim C_{3,3}^2-\dim C_{3,3}^3=8-4=4
.}
Since $\fr{g}\in\mathcal{M}$, we have $\dim \fr{g}_{-3,-3}\geq 4$. This shows that $\dim\fr{g}_{\bb{C}}\geq 10$. Therefore $\dim\fr{g}_{\bb{C}}\geq 8$. 
\end{proof}
We second show that the following. The following is the case of $rk(\pi_1(X,x))\leq 6$.
\begin{thm}\label{rk<=6}
Let $X$ be a smooth quasi-projective variety. If $\pi_1(X,x)$ is torsion-free nilpotent and $rk(\pi_1(X,x))\leq 6$, then $\pi_1(X,x)$ is a lattice in one of the following simply-connected nilpotent Lie groups:%すべてのlatticeではない
\ali{
\bullet\ &\bb{R}&\text{ if }b_1(X)=1,\\
\bullet\ &\bb{R}^2\text{ or }H_3(\bb{R})&\text{ if }b_1(X)=2,\\
\bullet\ &\bb{R}^3\text{ or }\bb{R}\times H_3(\bb{R})&\text{ if }b_1(X)=3,\\
\bullet\ &\bb{R}^4, \bb{R}^2\times H_3(\bb{R}), H_3(\bb{R})\times H_3(\bb{R})\text{ or }H_5(\bb{R})&\text{ if }b_1(X)=4,\\
\bullet\ &\bb{R}^5, \bb{R}^3\times H_3(\bb{R})\text{ or }\bb{R}\times H_5(\bb{R})&\text{ if }b_1(X)=5,\\
\bullet\ &\bb{R}^6&\text{ if }b_1(X)=6 
}  
where $H_{2k+1}(\bb{R})$ is $(2k+1)$-dimensional Heisenberg group.
\end{thm}
\begin{proof}
By \cref{twogen} and \cref{threegen}, we see that \cref{rk<=6} holds if $b_1(X)\leq 3$. If $b_1(X)=\dim H^1(\fr{g}_{\bb{C}})\geq 4$ and the step length of $\fr{g}_{\bb{C}}$ is greater than or equal to $4$, then we have 
\ali{
6&\geq rk(\pi_1(X,x))\\
&= \dim \fr{g}_{\bb{C}}\\
&=\dim\bo_{i\geq 1}C^i\fr{g}_{\bb{C}}/C^{i+1}\fr{g}_{\bb{C}}\\
&=\dim C^1 \fr{g}_{\bb{C}}/C^{2}\fr{g}_{\bb{C}}+\dim C^2\fr{g}_{\bb{C}}/C^{3}\fr{g}_{\bb{C}}+\dim C^3\fr{g}_{\bb{C}}/C^{4}\fr{g}_{\bb{C}}+\dim C^4\fr{g}_{\bb{C}}/C^{5}\fr{g}_{\bb{C}}+\cdots\\
&=\dim H^1(\fr{g}_{\bb{C}})+\dim C^2\fr{g}_{\bb{C}}/C^{3}\fr{g}_{\bb{C}}+\dim C^3\fr{g}_{\bb{C}}/C^{4}\fr{g}_{\bb{C}}+\dim C^4\fr{g}_{\bb{C}}/C^{5}\fr{g}_{\bb{C}}+\cdots\\
&\geq 4+1+1+1+\cdots\geq 7.
}
This shows that if $b_1(X)\geq 4$, then the step length of $\fr{g}_{\bb{C}}$ is $1$ (abelian) or $2$ or $3$. We consider the case of $b_1(X)=4$. By \cref{geq8}, we see that $\dim\fr{g}_{\bb{C}}\geq 8$ if $\fr{g}_{\bb{C}}$ is $3$-step. Thus $\fr{g}_{\bb{C}}$ is abelian or $2$-step nilpotent. Therefore, by the classification of complex nilpotent Lie algebras (\cite{Nil}), $\pi_1(X,x)$ is isomorphic to a lattice in one of the following Lie groups:
\ali{
\bb{R}^4, \bb{R}^2\times H_3(\bb{R}), H_5(\bb{R})\text{ or }H_3(\bb{R})\times H_3(\bb{R})
}
where $H_{2k+1}(\bb{R})$ is $(2k+1)$-dimensional Heisenberg group. We consider the case of $b_1(X)=5$. By the same argument as in the case of $b_1(X)=4$, we can show that $\fr{g}_{\bb{C}}$ is abelian or $2$-step nilpotent. Thus, by the classification of complex nilpotent Lie algebras (\cite{Nil}), $\pi_1(X,x)$ is isomorphic to a lattice in one of the following Lie groups:
\ali{
\bb{R}^5, \bb{R}^3\times H_3(\bb{R})\text{ or }\bb{R}\times H_5(\bb{R}).
}
It is clear that $\pi_1(X,x)$ is isomorphic to a lattice in $\bb{R}^6$ if $b_1(X)=6$. 
\end{proof}
Finally, we show the following theorem. The following is the case of $rk(\pi_1(X,x))=7$. 
\begin{thm}\label{rk=7}
Let $X$ be a smooth quasi-projective variety. If $\pi_1(X,x)$ is torsion-free nilpotent and $rk(\pi_1(X,x))=7$, then $\pi_1(X,x)$ is isomorphic to a lattice in a simply-connected real  nilpotent Lie group such that the complexification $\fr{n}_{\bb{C}}$ of the Lie algebra $\fr{n}$ of $N$ is isomorphic to one of the following Lie algebras: 
\ali{
\bullet\ &\bb{C}^7&\text{ if }b_1(X)=7,\\ 
\bullet\ &\bb{C}^4\o\fr{n}_3(\bb{C}), \bb{C}^2\o L_{5,4} \text{ or }\fr{n}_7^{152}&\text{ if }b_1(X)=6,\\
\bullet\ &\bb{C}\o\fr{n}_3(\bb{C})\o\fr{n}_3(\bb{C})\text{ or }(\text{ other $2$-step nilpotent such that $\dim H^1(\fr{n})=5$})&\text{ if }b_1(X)=5,\\
\bullet\ &\fr{n}_7^{142}\text{ or }\fr{n}_7^{143}&\text{ if }b_1(X)=4
}
where $L_{5,4}$ is the complexification of the Lie algebra of $H_5(\bb{R})$, $\fr{n}_7^{152}$ is the complexification of the Lie algebra of $H_7(\bb{R})$, $\fr{n}_7^{142}$ and $\fr{n}_7^{143}$ are given by
\ali{
\fr{n}_7^{142}=\ang{X_1,\dots,X_7}: [X_1,X_i]=X_{i-1}\text{ for i=3,5,7 },[X_3,X_5]=X_4,[X_5,X_7]=X_2,
}
and
\ali{
\fr{n}_7^{143}=\ang{X_1,\dots,X_7}: [X_1,X_i]=X_{i-1}\text{ for i=3,5,7 },[X_3,X_5]=X_6,[X_5,X_7]=X_2
.}
\end{thm}
\begin{proof}
By \cref{twogen} and \cref{threegen}, we see that $rk(\pi_1(X,x))\leq 4$ if $b_1(X)\leq 3$. Thus $b_1(X)\geq 4$. Let $N$ be a simply-connected nilpotent Lie group such that $\pi_1(X,x)$ is isomorphic to a lattice in $N$ and $\fr{g}$ the Lie algebra of $N$. By \cref{geq8} and the same argument as in \cref{rk<=6}, we see that $\fr{g}_{\bb{C}}$ is abelian or $2$-step nilpotent. We first assume that $b_1(X)=\dim H^1(\fr{g})=4$ and $\fr{g}_{\bb{C}}$ is $2$-step nilpotent. We define 
\ali{
(x,y,z):=(\dim H_{1,0}^1,\dim H_{0,1}^1,\dim H_{1,1}^1).
}
\begin{claim}
$(x,y,z)=(2,2,0)$.
\end{claim}
\begin{proof}
Assume that $(x,y,z)=(1,1,2)$. Let $X_1,\overline{X_1}$ and $Y_1,\overline{Y_1}$ be elements of a basis such that the dual elements $x_1,\overline{x_1},y_1,\overline{y_1}$ form a basis $H^1(\fr{g})=H_{1,0}^1\o H_{0,1}^1\o H_{1,1}^1=\ang{[x_1]}_{1,0}\o\ang{[\overline{x_1}]}_{0,1}\o\ang{[y_1],[\overline{y_1}]}_{1,1}$. We can show that $H_{3,1}^2,H_{1,3}^2\neq \0$ if $[X_1,\overline{X_1}]=0$. This contradicts the condition {\bf (W)}. Hence $[X_1,\overline{X_1}]\neq 0$. In addition, we can show that $H_{3,3}^2\neq \0$. Moreover we can show that $H_{3,3}^2\neq \0$ if $(x,y,z)=(0,0,4)$ by the same argument as in the case $[X_1,\overline{X_1}]=0$. Thus $(x,y,z)=(2,2,0)$.
\end{proof}
Let $\fr{g}_{-1,0}=\ang{X_1,X_2}$ and $\fr{g}_{0,-1}=\ang{\overline{X_1},\overline{X_2}}$. Since $\dim\fr{g}_{\bb{C}}=7$, we have $\dim\fr{g}_{-1,-1}=3$.
 By the classification of nilpotent Lie algebras of dimension up to $7$ over $\bb{C}$ (\cite{Nil}), we obtain the following lemmas. In particular, the following immediately follows from the classification of nilpotent Lie algebras over $\bb{C}$ in \cite{Nil}.
\begin{claim}\label{rk7b=4}
$\fr{g}_{\bb{C}}$ is isomorphic to one of the following Lie algebra:
\ali{
\fr{n}_7^{142}=\ang{X_1,\dots,X_7}: [X_1,X_i]=X_{i-1}\text{ for i=3,5,7 },[X_3,X_5]=X_4,[X_5,X_7]=X_2
}
\ali{
\fr{n}_7^{143}=\ang{X_1,\dots,X_7}: [X_1,X_i]=X_{i-1}\text{ for i=3,5,7 },[X_3,X_5]=X_6,[X_5,X_7]=X_2
.}
\end{claim}
\begin{proof}
We may assume that $Z_1:=[X_1,\overline{X_1}],Z_2:=[X_1,\overline{X_2}]$ and $Z_3:=[X_2,\overline{X_1}]$ form a basis of $\fr{g}_{-1,-1}$. Let $[X_2,\overline{X_2}]=aZ_1+bZ_2+cZ_3$. Note that
\ali{\overline{Z_1}&=\overline{[X_1,\overline{X_1}]},=[\overline{X_1},X_1]=-Z_1,\\
\overline{Z_2}&=\overline{[X_1,\overline{X_2}]}=[\overline{X_1},X_2]=-Z_3,\\
\overline{Z_3}&=-Z_2.
}
Let $e_1:=X_1,e_2:=\overline{X_1},e_3:=\overline{X_2},e_4:=X_2,e_5:=Z_1,e_6:=Z_2,e_7:=Z_3$. Then we have the structure constants $\{ C_{ij}^k\}_{i<j}$ such that
\ali{
C_{ij}^k=
\begin{cases}
1&(i,j,k)=(1,2,5),(1,3,6),(4,2,7)\\
a&(i,j,k)=(4,3,5)\\
b&(i,j,k)=(4,3,6)\\
c&(i,j,k)=(4,3,7)\\
0&\text{ otherwise}.
\end{cases}
}
By the classification of complex nilpotent Lie algebras (\cite{Nil}) and since the dimension of the center of $\fr{g}_{\bb{C}}$ is $3$, we have
\ali{
\fr{g}_{\bb{C}}\cong \fr{n}_7^{142}\text{ or }\fr{n}_7^{143}\text{ or }\fr{n}_7^{144}\text{ or }\fr{n}_7^{145}.
}
By the direct computation, we have
\ali{
(\dim H^2(\fr{n}_7^i),\dim H^3(\fr{n}_7^i))
=
\begin{cases}
(11,14)&i=142\\
(11,16)&i=143\\
(11,17)&i=144\\
(12,18)&i=145.
\end{cases}
} 
Since
\ali{
(\dim H^2(\fr{g}_{\bb{C}}),\dim H^3(\fr{g}_{\bb{C}}))
=
\begin{cases}
(11,14)&a\neq 0\text{ or }b\neq 0\\
(11,16)&a=b=0,\\
\end{cases}
}
we have
\ali{
\fr{g}_{\bb{C}}\cong \fr{n}_7^{142}\text{ or }\fr{n}_7^{143}
.}
\end{proof}  
Note that $\fr{g}_{\bb{C}}$ is $2$-step or abelian if $b_1(X)=\dim H^1(\fr{g}_{\bb{C}})=5,6$ or $7$. By the classification of complex nilpotent Lie algebras (\cite{Nil}), we have that $\fr{g}_{\bb{C}}$ is isomorphic to one of the following Lie algebras:
\ali{
&\bb{C}^7&\text{ if }b_1(X)=7,\\
&\bb{C}^4\o\fr{n}_3(\bb{C}), \bb{C}^2\o L_{5,4} \text{ or }\fr{n}_7^{152}&\text{ if }b_1(X)=6,\\
&\bb{C}\o\fr{n}_3(\bb{C})\o\fr{n}_3(\bb{C})\text{ or }(\text{other $2$-step nilpotent such that $\dim H^1(\fr{n})=5$})&\text{ if }b_1(X)=5.
}
\end{proof}
\section{The case of $rk(\pi_1(X,x))=8$}
Let $X$ be a smooth quasi-projective algebraic variety. Assume that $\pi_1(X,x)$ is torsion-free and $3$-step nilpotent. Let $N$ be the simply-connected nilpotent Lie group such that $\pi_1(X,x)$ is a lattice in $N$. By \cref{geq8}, we have $rk(\pi_1(X,x))=\dim N=\dim \fr{g}\geq 8$ where $\fr{g}$ is the Lie algebra of $N$. In this section, we consider the case of $rk(\pi_1(X,x))=8$. We give an example of $3$-step nilpotent Lie group whose lattice may be isomorphic to the fundamental group of a smooth quasi-projective variety.  
\begin{prop}\label{8dim}
Let $\fr{g}=\ang{X_1,X_2,Y_1,Y_2,Z_1,Z_2,A,B}$ such that 
\ali{
[X_1,Y_1]=Z_1=[X_2,Y_2], [X_2,Y_1]=Z_2=[X_1,Y_2],\\
[X_1,Z_1]=A=[X_2,Z_2],[Y_1,Z_1]=B=[Y_2,Z_2]
.}
Then $\fr{g}$ is $3$-step and $\fr{g}\in \mathcal{M}$.
\end{prop}
\begin{proof}
We can easily check that $\fr{g}$ is $3$-step. We show that $\fr{g}\in\mathcal{M}.$ We define the bigrading on $\fr{g}_{\bb{C}}$ to be 
\ali{
\fr{g}_{\bb{C}}&=\ang{A_1,A_2}_{-1,0}\o\ang{\overline{A_1},\overline{A_2}}_{0,-1}\o\ang{B_1,B_2}_{-1,-1}\o\ang{C_1}_{-2,-1}\o\ang{\overline{C_1}}_{-1,-2}
}
where $A_1:=X_1+iY_1,A_2:=X_2+iY_2,B_1:=-2iZ_1,B_2:=-2iZ_2$ and $C_1:=-2i(A+iB)$. Thus the bigrading defines a mixed Hodge structure on $\fr{g}$.
Let $a_1,a_1,\overline{a_1},\overline{a_2},b_1,b_2,c_1$ and $\overline{c_1}$ be the dual basis. Since 
\ali{
[A_1,\overline{A_1}]=B_1,[A_2,\overline{A_2}]=B_2,\\
[A_1,B_1]=C_1=[A_2,B_2],\\
[\overline{A_1},\overline{B_1}]=\overline{C_1}=[\overline{A_2},\overline{B_2}],
}
and $[X,Y]=0\text{ for other }X,Y\in \{A_1,\dots, \overline{C_1}\},$ we have 
\ali{
da_1&=0=da_2=d\overline{a_1}=d\overline{a_2},\\
db_1&=-a_1\w \overline{a_1},db_2=-a_2\w \overline{a_2},\\
dc_1&=-a_1\w b_1-a_2\w b_2,d\overline{c_1}=\overline{a_1}\w b_1+\overline{a_2}\w b_2.
}
Hence the induced bigradings on the first and second cohomologies are given by
\ali{
H^1(\fr{g})=\ang{[a_1],[a_2]}_{1,0}\o\ang{[\overline{a_1}],[\overline{a_2}]}_{0,1}
}
and
\ali{
H^2(\fr{g})&=\ang{[a_1\w a_2]}_{2,0}\o\ang{[\overline{a_1}\w \overline{a_2}]}_{0,2}\o\ang{[a_1\w \overline{a_2}],[\overline{a_1}\w a_2]}_{1,1}\\
&\o\ang{[a_1\w b_1]}_{2,1}\o\ang{[\overline{a_1}\w b_1]}_{1,2}.
}
Therefore $\fr{g}\in\mathcal{M}.$
\end{proof}
\begin{rem}
Let $\fr{g}=\ang{X_1,X_2,Y_1,Y_2,Z_1,Z_2,A,B}$ be the Lie algebra in \cref{8dim}. Then $\fr{g}$ is isomorphic to the Lie subalgebra $\overline{\fr{g}}$ of $\fr{gl}_9(\bb{R})$ given by
\ali{
\overline{\fr{g}}:=\left \{\begin{pmatrix}
0&0&y_2&y_1&-z_2&-z_1&0&0&3b\\
0&0&x_2&x_1&0&0&-z_2&-z_1&3a\\
0&0&0&0&x_1&x_2&-y_1&-y_2&2z_2\\
0&0&0&0&x_2&x_1&-y_2&-y_1&2z_1\\
0&0&0&0&0&0&0&0&y_2\\
0&0&0&0&0&0&0&0&y_1\\
0&0&0&0&0&0&0&0&x_2\\
0&0&0&0&0&0&0&0&x_1\\
0&0&0&0&0&0&0&0&0\\
\end{pmatrix}
;
x_1,x_2,y_1,y_2,z_1,z_2,a,b\in\bb{R}\right \}
,}
which can be proved by constructing the isomorphism $\phi :\fr{g}\rightarrow \overline{\fr{g}}$ such that
\ali{
\phi (X_1)&:=e_{2,4}+e_{3,5}+e_{4,6}+e_{8,9}\\
\phi (X_2)&:=e_{2,3}+e_{3,6}+e_{4,5}+e_{7,9}\\
\phi (Y_1)&:=e_{1,4}-e_{3,7}-e_{4,8}+e_{6,9}\\
\phi (Y_2)&:=e_{1,3}-e_{3,8}-e_{4,7}+e_{5,9}\\
\phi (Z_1)&:=-e_{1,6}-e_{2,8}+2e_{4,9}\\
\phi (Z_2)&:=-e_{1,5}-e_{2,7}+2e_{3,9}\\
\phi (A)&:=3e_{2,9}\\
\phi (B)&:=3e_{1,9}.
}
Furthermore, the corresponding simply-connected nilpotent Lie group $N(\bb{R})$ is given by
\ali{
N(\bb{R})=\left \{\begin{pmatrix}
1&0&y_2&y_1&-z_2&-z_1&0&\frac{1}{2}(y_1+y_2)&3b\\
0&1&x_2&x_1&0&\frac{1}{2}(x_1+x_2)&-z_2&-z_1&3a\\
0&0&1&0&x_1&x_2&-y_1&-y_2&2z_2\\
0&0&0&1&x_2&x_1&-y_2&-y_1&2z_1\\
0&0&0&0&1&0&0&0&y_2\\
0&0&0&0&0&1&0&0&y_1\\
0&0&0&0&0&0&1&0&x_2\\
0&0&0&0&0&0&0&1&x_1\\
0&0&0&0&0&0&0&0&1\\
\end{pmatrix}
;
x_1,x_2,y_1,y_2,z_1,z_2,a,b\in\bb{R}\right \}
.}
\end{rem}
\begin{prob}
Does there exist a smooth quasi-projective variety $M$ such that $\pi_1(M,x)$ is  a lattice in $N(\bb{R})$?
\end{prob}
\section{An algebraic remark}%gradingとbigradingの差を強調する表はいらなくて反例はリー代数の表示だけでいい。6次元までは
\subsection{Gap between gradings and bigradings}
In \cite{grading}, we study lattices in simply-connected nilpotent Lie groups which could be realized as the fundamental groups of smooth quasi-projective varieties by using graded Lie algebras satisfying certain conditions. The following, the conditions of graded Lie algebras, immediately follows from \cref{W}.
\begin{cor}[\cite{Morgan,grading}]
Let $M$ be a any smooth quasi-projective variety, $x\in M$ based point.
 We assume that  the fundamental group $\pi_{1}(M,x)$ is a lattice in a simply connected nilpotent Lie group $N$.
 Then there exists a grading $\fr{n}_{\bb{C}}=\dis{\bo_{i\leq -1}\fr{n}_i}$ such that:
\\\ \indent
{\bf (W):} For the induced grading $H^*(\fr{n}_{\bb{C}})=\dis{\bo_{i\geq 1}H_i^*}$, we have
\ali{H^1(\fr{n}_{\bb{C}})=H_1^1\o H_2^1}
\\\ \indent
and
\ali{H^2(\fr{n}_{\bb{C}})=H_2^2\o H_3^2\o H_4^2.} 
\\\ \indent
{\bf (H):} For any odd integer $k$, $\dim{\fr{n}_k}$ and $\dim{H_k^j}$ are even.
\end{cor}
The condition {\bf (W)} holds by taking the total grading $\fr{n}_i:=\bo_{p+q=i}\fr{n}_{p,q}$. The condition {\bf (H)} holds by the Hodge symmetry $\overline{\fr{n}_{p,q}}\equiv \fr{n}_{q,p}\mod \bo_{s+t<p+q}\fr{n}_{s,t}$.
In this paper, we obtain sharper criteria for a lattice in a simply-connected nilpotent Lie group to be the fundamental group of a smooth quasi-projective variety. Furthermore, our classification in dimensions up to 6 includes the results of \cite{grading}.
\begin{rem}
The following Lie algebras
\ali{
L_{5,9}=\ang{X_1,X_2,X_3,X_4,X_5}:&[X_1,X_2]=X_3,[X_2,X_3]=X_4,[X_1,X_3]=X_5,\\
L_{6,9}=L_{5,9}\o\bb{C},\hspace{2.1cm}&\\
L_{6,21}(-1)=\ang{X_1,X_2,X_3,X_4,X_5,X_6}:&[X_1,X_2]=X_3, [X_2,X_3]=X_4,[X_1,X_3]=X_5,\\&[X_1,X_4]=X_6=[X_2,X_5],\\
L_{6,22}(0)=\ang{X_1,X_2,X_3,X_4,X_5,X_6}:&[X_2,X_4]=X_5, [X_4,X_1]=X_6=[X_2,X_3],\\
L_{6,24}(0)=\ang{X_1,X_2,X_3,X_4,X_5,X_6}:&[X_1,X_3]=X_4, [X_3,X_4]=X_5,[X_1,X_4]=X_6=[X_3,X_2]
}
and
\ali{
L_{6,24}(1)=\ang{X_1,X_2,X_3,X_4,X_5,X_6}:&[X_1,X_2]=X_3, [X_2,X_3]=X_5=[X_2,X_4],[X_1,X_3]=X_6
}
do not admit a bigrading satisfying {\bf (W)}. However, these admit  gradings satisfying {\bf (W)} and {\bf (H)} $($ see \cite{grading}$)$. Thus the bigradings provide sharper criteria for lattices in simply-connected nilpotent Lie groups to be realized as the fundamental groups of smooth quasi-projective varieties than gradings.
\end{rem}

\end{document}